OpenStudy (anonymous):

In order for a satellite to remain in a circular orbit above the same spot on earth the satellite must be 35,800 km above the equator. Write an equation for the orbit of the satellite. Use the center of Earth as the origin and 6400 km for the radius of the Earth.

OpenStudy (primeralph):

in terms of energy or potential or what?

OpenStudy (shamim):

\[\frac{ mv ^{2} }{ R+h }=\frac{ GMm }{ (R+h)^{2} }\]

OpenStudy (shamim):

i think its a question of physics